Ace Industrial Machines issued 115,000 zero coupon bonds 7 years ago. The bonds originally had 30 years to maturity with a yield to maturity of 6 percent. Interest rates have recently decreased, and the bonds now have a yield to maturity of 5.1 percent. The bonds have a par value of $2,000. If the company has a $76.2 million market value of equity, what weight should it use for debt when calculating the cost of capital? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., .1616.)
Answer #1

par value =   2000
Years to maturity= 30-7=   23
  
Yield to Maturity = 5.1% or   0.051
  
Price of Zero coupon bond formula = Face Value /(1+ interest or yield rate)^years to maturity  
2000/(1+0.051)^23  
637.0412091  
  
So, price of 115000 Zero coupon bond is 115000*637.0412091 =$   73259739.05


Value of Equity=   76200000


Weight of debt = Debt/(Debt + Equity)  
73259739.05   /(73259739.05+76200000)
0.4901637024  
  
So, weight of debt used for cost of capital is   0.4902

Please thumbs up